OpenStudy (anonymous):

How can I replace the underscores with numbers so the system has only one solution?: y=4x+7 _y=_x+_

OpenStudy (anonymous):

In order for the system to have only one solution, the lines given by the equations must intersect at only one point. Two examples would be: y=4x+7 and y=-4x+7 as well as y=4x+7 and -y=4x+7

OpenStudy (anonymous):

How do you know these two equations only intersect at only point? It is if the lines have the same x or y intercept, but the oposite sign?

OpenStudy (anonymous):

Those were just two examples. What I did, though, was replace "x" with "-x." Doing so reflects the graph of y=4x+7 over the y-axis.

OpenStudy (anonymous):

Thanks!